the length of a rectangle is 2 more than 3 times the width.if the perimeter is 92 find the dimensions of the rectangle

Answer by macston(5194)

.
W=width; L=length=3W+2
.
P=2(L+W)
92=2(3W+2+W)
46=4W+2
44=4W
11=W
.
L=3W+2=3(11)+2=33+2=35
ANSWER: The rectangle is 35 x 11.
.
CHECK:
P=2(L+W)
92=2(35+11)
92=2(46)
92=92